# Degree Matrix Comparison for Graph Alignment

### v2

@@ Abstract
Graph alignment considers the optimal node correspondence across networks. To advance unsupervised graph alignment algorithms on plain graphs, we propose Degree Matrix Comparison (DMC). Through extensive experiments and mathematical motivations, we demonstrate the potential of this method. Remarkably, DMC achieves up to $99\%$ correct node alignment for $90\%$-overlap graphs and $100\%$ accuracy for isomorphic graphs. Additionally, we propose a reduced version of DMC (Greedy DMC) that provides a solution to the graph alignment problem with lower time complexity. DMC could significantly impact graph alignment, offering a reliable solution for the task.
@@ 1 Introduction
Graph alignment is a critical task with applications across various domains, like social networks, biology, and cybersecurity. It is also highly related to the subgraph isomorphism problem in mathematics and computer science. Hence, improving graph alignment methods holds significant value. Supervised or semi-supervised methods typically exploit node attributes, such as semantic features of users [1, 2], but these methods break down when users disguise their identities [3]. WAlign is a recent unsupervised method, but also harnesses the power of attributes [4]. We focus on the unsupervised graph alignment problem for unattributed plain graphs, leveraging geometric properties.
Our proposed method, Degree Matrix Comparison (DMC), is a method that fits the task. Unsupervised methods with a focus on graph geometry include: REGAL [5], FINAL [6], Klau [7], and IsoRank [8]. For a more comprehensive view of past unsupervised alignment methods, see [9]. DMC was constructed with two core ideas in mind: degree is the most accurate and direct description of local structure; moreover, in representing geometric structure, it makes sense to include both local and global information, like in [10, 11], especially in compact matrix form (adjacency matrices and Laplacian matrices).
We propose foregoing the need to know exactly which nodes are connected, like in an adjacency matrix, and record accurately the degree of a node and its neighbors’ degrees in a matrix. We show that DMC works well for aligning real world graphs with high clustering and link complexity.
@@ 2 Problem Set Up
Geometrically, $G$ could represent an object with vertices and edges (links) that connect some of the vertices. The presence of $\{v_{1},v_{2}\}\in E$ indicates there is an edge between the vertices $v_{1},v_{2}$. When we mention graphs in this article, we typically talk about them in a manner that assumes a geometric interpretation. But there are other ways to interpret what we call a “graph”; a most direct example is the adjacency matrix, which is a binary matrix representation of graphs.
Definition 2.2 (Graph Alignment).
@@
To perform graph alignment, we generate a pair of graphs with common structure. We introduce two graph sampling methods–these methods sample size-wise manageable smaller graphs from larger complex networks.
@@ 2.1 Random Walk Graph Sampling
One of our methods utilizes random walk on networks to extract subgraphs [3, 12].
Definition 2.3 (Random Walk (on graphs)).
@@
To construct $G_{1}=(V_{1},E_{1})\subseteq G_{s}$ and $G_{2}=(V_{2},E_{2})\subseteq G_{s}$ for graph alignment, we perform random walk on $G_{s}$ to first obtain a set of nodes that serves as the common set of nodes $C=V_{1}\bigcap V_{2}\subset G_{s}$. The number of vertices we want to stop at, for the random walk, is $|C|=np$, where $n$ is like before the number of distinct nodes in $G_{s}$ and $p$ is what we call the “overlap percentage” (with respect to $G_{s}$). We then take arbitrary $H_{1},H_{2}\subseteq(V_{s}\setminus C)$ such that $|H_{1}|=|H_{2}|$ and $V_{s}\setminus C=H_{1}\sqcup H_{2}$ (disjoint union). Let $V_{1}=C\sqcup H_{1}$ and $V_{2}=C\sqcup H_{2}$, then $|V_{1}|=|V_{2}|$. $E_{1}\subseteq E_{s}$ is the inherited set of edges between nodes in $V_{1}$ from $E_{s}$, likewise for $E_{2}$.
@@
Figure 1 is a demonstration for what we are trying to take off the entire graph. There are three green circles containing different regions of the graph; the two larger ones contain the same number of nodes, and the smallest circle contains the common region between the two possible sampled graphs through random walk. Note that this is not the only sampling result possible.
@@ 2.2 Edge Deletion Graph Sampling
The second graph sampling method is learned from [5]. Suppose $G_{s}$ is the same graph as in section 2.1, then let $G_{1}=G_{s}=(V_{1},E_{1})$, and we delete edges of $G_{1}$ with probability $p_{d}$ to obtain another graph, which is $G_{2}$. If $G_{2}=(V_{2},E_{2})$, then $V_{2}=V_{1}$, and $E_{2}\subseteq E_{1}$.
@@ 3.1 Degree Matrix Comparison
Definition 3.1 (Degree Matrices).
@@
The order of rows in $M_{1}$ and $M_{2}$ is arbitrary. After obtaining $M_{1}$ and $M_{2}$, we use the Munkres assignment algorithm to minimize cost between rows of the matrices, producing a map between rows of $M_{1}$ and $M_{2}$. We migrate this assignment directly to the assignment $f:V_{1}\rightarrow V_{2}$ between nodes. We call this entire process the Degree Matrix Comparison (DMC).
@@
For hands-on examples to demonstrate the concept of a degree matrix and the DMC process, see Appendix A. If the reader is not familiar with the Munkres assignment algorithm, see Appendix B for details. In practice (implementing in code), the Munkres assignment algorithm can be replaced by a computationally more efficient linear_sum_assignment function from scipy.optimize in Python which yields the same assignment as Munkres [13]. But in terms of comprehension, the Munkres assignment algorithm is more straightforward.
The complexity of DMC (either theoretically with Munkres or computationally with linear_sum_assignment) is at most $O(N^{3})$ for a pair of graphs that each has $N$ nodes [14, 15]. When $m<<N$, the complexity is $O(N^{2}\cdot m)<<O(N^{3})$.
3.2 Variation: Greedy DMC
@@
The complexity of DMC can be reduced through replacing direct Munkres assignment with a combination of the auction algorithm and Munkres assignment algorithm, but accuracy is compromised [16]. The auction algorithm, with $O(N^{2})$ complexity and assignment threshold $\epsilon$, serves as a preparation step to reduce reliance on the Munkres assignment algorithm. We call this reduced version the Greedy DMC.
@@
For details about the auction algorithm, one could refer to Appendix C. Here we bring to attention the case where $\epsilon=0$. While this is a special case of the Greedy DMC algorithms with different $\epsilon$ thresholds, it is not “Greedy” in any sense. In fact, this is an algorithm that first finds exact matches (same number of first neighbors and its neighbors’ degrees follow exact same distribution) between rows in $M_{1}$ and $M_{2}$, then performs the Munkres assignment algorithm on other rows. As can be seen in section 5 with the results, Greedy DMC’s $\epsilon=0,10$ cases are very effective algorithms like DMC, and are worth studying in the future.
@@ 4 Theoretical Motivations
One theoretical motivation for the method is that degree matrices remain invariant in response to arbitrary node labeling. Moreover, a degree matrix ($N\times m$) encodes information more compactly than an adjacency matrix ($N\times N$). Additionally, combining local (degrees) and global (assignment) perspectives is important because the Friendship paradox shows how local observations can be completely shifted when we examine the same neighborhood globally [17, 18].
The Friendship paradox, in general terms, is the observation that friends of a person typically have more friends than that person. There are many ways to mathematically formulate the problem, and this leads to a whole other field of study; we will not expand on the details here.
Intuitively, our method works well for heterogeneous graphs for the following reason. Suppose there is a graph that has nodes with degrees that are almost all unique. (We phrase it this way with the following consideration in mind: If we allow at most one edge between every pair of vertices on a graph $G_{example}$ with $n_{example}$ vertices and require at least degree one for all nodes, then the maximum degree is at most $n_{example}-1$ for every node, and in this case it is impossible to create a bijection between possible degrees and the nodes.) Then most nodes have row vectors that are relatively unique and distinguishable, compared to, for example, a large set of degree one nodes all connected to a common center.
5 Experimental Results and Analysis
5.1 Real World Networks
Here is a table that concludes the wide range of datasets we worked with.
COSNET [19] Flickr; Last.fm; Myspace
SNAP [20] Facebook; YouTube; CA, PA Roadmaps; PPI
NetworkRepository [21] All other networks.
Method Correct Percentage Approx. Complexity
DMC $96.09\%$ $\sim O(n^{2})-O(n^{3})$
Greedy DMC ($\epsilon=100$) $50.46\%$ $\sim O(n^{2})-O(n^{3})$
REGAL [5] $80\%<p<90\%$ $\sim O(n)-O(n^{2})$
FINAL [6] $35\%<p<45\%$ $\sim O(n^{2})$
Klau [7] $20\%<p<30\%$ $\sim O(n^{5})$
IsoRank [8] $5\%<p<15\%$ $\sim O(n^{4})$
Two-Step $4.42\%$ $\sim O(n^{2})-O(n^{3})$
Euc Dist. $0.15\%$ (near zero) $\sim O(n^{2})-O(n^{3})$

### v3

@@ Abstract
The graph alignment problem, which considers the optimal node correspondence across networks, has recently gained significant attention due to its wide applications. There are graph alignment methods suited for various network types, but we focus on the unsupervised geometric alignment algorithms. We propose Degree Matrix Comparison (DMC), a very simple degree-based method that has shown to be effective for heterogeneous networks. Through extensive experiments and mathematical proofs, we demonstrate the potential of this method. Remarkably, DMC achieves up to 99% correct node alignment for 90%-overlap networks and 100% accuracy for isomorphic graphs. Additionally, we propose a reduced Greedy DMC with lower time complexity and Weighted DMC that has demonstrated potential for aligning weighted graphs. Positive results from applying Greedy DMC and the Weighted DMC furthermore speaks to the validity and potential of the DMC. The sequence of DMC methods could significantly impact graph alignment, offering reliable solutions for the task.
[…]
123
@@ 1 Introduction
Graph alignment is a critical task with applications across various domains, like social networks, biology, and cybersecurity. In social networks, graph alignment can identify the same user across multiple platforms, providing insights into user behavior and preferences. In biology, it helps in comparing molecular structures and understanding functional similarities between biological networks. In cybersecurity, aligning graphs can detect coordinated attacks and identify malicious entities by correlating data from different sources. Graph alignment is also highly related to the subgraph isomorphism problem in mathematics and computer science, which is an NP-complete problem. Improving the efficiency and simplifying implementation steps of graph alignment algorithms are important in artificial intelligence because they help us learn patterns on networks.
There are supervised graph alignment methods that work well already [18]. This is not surprising as supervised and semi-supervised methods exploit node attributes, such as semantic features of users [25, 30]. However, they break down when users disguise their identities and node attributes are no longer reflective of local properties [11]. Moreover, they require some knowledge of the node labeling on a network prior to the alignment. There are also unsupervised methods that rely on the power of node attributes, like WAlign [10]. We, however, focus on the unsupervised graph alignment problem for unattributed plain graphs, mainly leveraging geometric properties.
Our proposed method, Degree Matrix Comparison (DMC), is a method that fits the task. We list some previous unsupervised methods that exploit graph geometries: REGAL, which is a representation learning-based method [12]; FINAL, which mainly utilizes graph topology but enhances it with node attributes [31]; Klau, which utilizes a Lagrangian relaxation approach in an optimization problem [13]; and IsoRank, which constructs an eigenvalue problem for every pair of input graphs and then uses k-partite matching techniques [23]. For a more comprehensive view of past unsupervised alignment methods, see Skitsas et al. [24].
DMC was constructed with the following two core ideas in mind. First of all, degree is the most accurate and direct description of local structure. Moreover, in representing geometric structure, it makes sense to include both local and global information, like in Kuchaiev et al. [14], Milano et al. [19], especially in compact matrix form (like in adjacency matrices and Laplacian matrices). This is not just because “the more the better”, rather it is a consequence of the Friendship Paradox. More specifically, we propose foregoing the need to know exact connections between nodes, like in an adjacency matrix, and record accurately two-layer exact information–degree of a node and its neighbors’ degrees in a matrix–in a clever way, for later use of the Hungarian algorithm. Since we are aligning unattributed graphs, a method that is based on information about links between labeled nodes is not rigorous, because the assignment of node IDs could be random on the pair of graphs to be aligned. Although the Hungarian algorithm is widely known and has been integrated into other graph alignment methods [4], our approach merits further attention and research due to its exceptional applicability to real heterogeneous networks–rather than restricted ideal graphs like k-partite graphs or complete graphs–with very simple implementation steps.
@@ 2 Problem Set Up
Geometrically, $G$ could represent an object with vertices and edges that connect some of the vertices. The presence of $\{v_{1},v_{2}\}\in E$ indicates there is an edge between the vertices $v_{1},v_{2}$. We talk about graphs in a manner that assumes a geometric interpretation. But there are other ways to interpret what we call a “graph”; a most direct example is the adjacency matrix, which is a binary matrix representation of graphs.
Definition 2 (Graph Alignment).
@@
To perform graph alignment, we generate a pair of graphs with common structure. We introduce two graph sampling methods that sample size-wise manageable smaller graphs from larger complex graphs.
@@ 2.1 Random Walk Graph Sampling
One of our methods utilizes random walk on graphs to extract subgraphs [11, 16].
Definition 3 (Random Walk (on graphs)).
@@
To construct $G_{1}=(V_{1},E_{1})\subseteq G_{s}$ and $G_{2}=(V_{2},E_{2})\subseteq G_{s}$ for graph alignment, we perform random walk on $G_{s}$ to first obtain a set of nodes that serves as the common set of nodes $C=V_{1}\bigcap V_{2}\subset G_{s}$. The number of vertices we want to stop at, for the random walk, is $|C|=np$, where $n$ is like before the number of distinct nodes in $G_{s}$ and $p$ is what we call the “overlap percentage" (with respect to $G_{s}$). We then take arbitrary $H_{1},H_{2}\subseteq(V_{s}\setminus C)$ such that $|H_{1}|=|H_{2}|$ and $V_{s}\setminus C=H_{1}\sqcup H_{2}$ (disjoint union). Let $V_{1}=C\sqcup H_{1}$ and $V_{2}=C\sqcup H_{2}$, then $|V_{1}|=|V_{2}|$. $E_{1}\subseteq E_{s}$ is the inherited set of edges between nodes in $V_{1}$ from $E_{s}$, likewise for $E_{2}$.
Figure 1 is a demonstration for what we are trying to take off the entire graph. There are three green circles containing different regions of the graph; the two larger ones contain the same number of nodes, and the smallest circle contains the common region between the two possible sampled graphs through random walk. This is not the only sampling possible.
@@ 2.2 Edge Deletion Graph Sampling
The second graph sampling method is from Heimann et al. [12]. Suppose $G_{s}$ is the same graph as in Section 2.1, then let $G_{1}=G_{s}=(V_{1},E_{1})$, and we delete edges of $G_{1}$ with probability $p_{d}$ to obtain another graph, which is $G_{2}$. If $G_{2}=(V_{2},E_{2})$, then $V_{2}=V_{1}$, and $E_{2}\subseteq E_{1}$.
@@ 3.1 Degree Matrix Comparison
Definition 4 (Degree Matrices).
@@
The order of rows in $M_{1}$ and $M_{2}$ is arbitrary. After obtaining $M_{1}$ and $M_{2}$, we use the Hungarian algorithm to minimize cost between rows of the matrices, producing a map between rows of $M_{1}$ and $M_{2}$. We migrate this assignment directly to the assignment $f:V_{1}\rightarrow V_{2}$ between nodes. We call this entire process the Degree Matrix Comparison (DMC).
We provide examples here to demonstrate the concept of a degree matrix and the DMC process. For the degree matrix formulation, let us examine the graph (assumed to be a local picture taken from a much larger graph) in Figure 3 with circled node as the origin. Then the boxed nodes are the first neighbors of the origin. Starting from the first neighbor in the upper left corner and going counter clock-wise, the degrees of the five first neighbors are $3,4,3,5,3$. Suppose (arbitrarily) $m$ in this case is $10$, then the row vector representation for the circled node would be $[3,3,3,4,5,0,0,0,0,0]$ in the degree matrix for the larger graph. We do this for every node in the larger graph which this smaller piece is taken from to build a degree matrix.
Figure 3: Example Graph to Demonstrate Degree Matrix Formulation
Figure 4: $F_{1}$
Figure 5: $F_{2}$
Now we demonstrate the DMC process. Consider the two graphs in Figures 4 and 5 as a pair that we are to align. Let the graph in Figure 4 be $F_{1}=(V_{F_{1}},E_{F_{1}})$ and the graph in Figure 5 be $F_{2}=(V_{F_{2}},E_{F_{2}})$. The node set $V_{l}$ of the largest overlapping subgraph consists of all five nodes, or: $V_{l}=V_{F_{1}}=V_{F_{2}}$. However, the graphs are not isomorphic, so the edge set $E_{l}$ of the overlapping subgraph is a strict subset of $E_{F_{1}},E_{F_{2}}$, which in mathematical notation is: $E_{l}\subsetneq E_{F_{1}},E_{F_{2}}$. Let the edge between node 3 and node 4 in $F_{2}$ be $w$. Then the overlapping subgraph takes the shape of $F_{2}\backslash\{w\}$. The degree matrices for $F_{1}$ and $F_{2}$ are
[…]
Then we use the Hungarian algorithm to minimize the cost of row alignment between the two matrices above. The algorithm aligns nodes $1,2,3,4,5$ of $F_{1}$ to nodes $2,5,1,3,4$ of $F_{2}$ (one could check understanding of the concepts with this result). This is not an optimal result, but it is expected since the graphs are not heterogeneous (the relationship between performance of DMC and heterogeneity of graph will be discussed in depth in Section 5). This example is purely for demonstrating DMC.
@@
In practice (implementing in code), the Hungarian algorithm can be replaced by a computationally more efficient linear_sum_assignment function from scipy.optimize in Python which yields the same assignment [6]. However, in terms of comprehension, the Hungarian algorithm is more straightforward. The complexity of DMC is at most $O(N^{3})$ for a pair of graphs that each has $N$ nodes [8, 27]. When $m<<N$, the complexity is $O(N^{2}\cdot m)<<O(N^{3})$.
3.2 Variation 1: Greedy DMC
@@
The complexity of DMC can be reduced through replacing the Hungarian algorithm with a combination of the auction algorithm and Hungarian algorithm, but accuracy is compromised [3]. The auction algorithm, with $O(N^{2})$ complexity and assignment threshold $\epsilon$, serves as a preparation step to reduce reliance on the Hungarian algorithm. We call this reduced version the Greedy DMC.
@@
We bring to attention the case where $\epsilon=0$. While this is a special case of the Greedy DMC algorithms with different $\epsilon$ thresholds, it is not “Greedy” in any sense. In fact, this is an algorithm that first finds exact matches between rows in $M_{1}$ and $M_{2}$, then performs the Hungarian algorithm on remaining rows. As can be seen in Section 5, Greedy DMC’s $\epsilon=0,10$ cases are very effective algorithms like DMC, and are worth studying in the future.
3.3 Variation 2: Weighted DMC
While the previous methods are meant for unweighted plain graphs, the Weighted DMC is designed for weighted graphs. Graphs can no longer be fully represented by just an adjacency matrix, which counts number of edges but makes no note of varying weights. The idea is simple: replace each entry in the degree matrix with its original value multiplied by the weight of the edge connecting the row’s node and the node the entry represents.
Weighted DMC performed well on multiple datasets, suggesting it deserves further research–much like Greedy DMC. In particular, its strong performance on protein-protein interaction (PPI) networks (often used to test graph alignment) indicates that DMC’s success is not coincidental.
@@ 4 Theoretical Motivations
4.1 Invariance
One theoretical motivation for the method is that degree matrices remain invariant in response to arbitrary node labeling. By invariant, we mean that degree matrices are in the same equivalence class defined by the following equivalence relation.
Theorem 1 (Equivalence Relation on Degree Matrices.).
We define a binary relation between two degree matrices: degree matrices $A,B$ are said to be related (i.e. $A\sim B$) if we can transform $A$ to become $B$ only through row swap operations. Moreover, we claim that this is an equivalence relation.
Proof.
A binary relation is an equivalence relation if and only if the relation is reflexive, symmetric, and transitive. For any matrix $A$, $A\sim A$ is true since we can obtain $A$ from performing no row swap operations on $A$, hence the relation is reflexive.
$A\sim B$ implies there is a series of row swaps that can transform $A$ to $B$. Therefore, we can also obtain $A$ from $B$ through reversing the row swap process and so $B\sim A$. Hence, the relation is symmetric.
$A\sim B$ and $B\sim C$ implies we can transform $A$ into $B$ through row swaps, and also transform $B$ into $C$ through row swaps. This implies we can obtain $C$ from $A$ through first performing row swaps that transform $A$ into $B$ and then applying more row swaps that transform $B$ into $C$. Hence we can obtain $C$ from $A$ through row swaps and $A\sim C$, making the binary relation transitive. ∎
This is crucial for graph alignment for unattributed graphs (or on attributed graphs but not using attributes).
4.2 Performance Analysis
We provide an analysis of performance of the DMC, which shows that the DMC is effective in differentiating high- and low-degree nodes. In a heterogeneous graph similar to one generated by the Barabási-Albert model (introduced in Section 5), a significant amount of nodes are either a concentrated hub center or close to a leaf due to the formation of hubs. Therefore, we can assume the row vector in a degree matrix for a high-degree node (a hub) looks like
$[1,1,1,1,1,1,1,2,3,0,0]$ (1)
just as an example. The main properties of this row vector are that entries are small numbers (low-degree nodes) and there are many non-zero entries, due to node’s large degree. For a leaf, the row vector will look like
$[100,0,0,0,0,0,0],$ (2)
for instance. In these low-degree node row vectors, we observe that most entries are zero, and the few non-zero entries tend to be large. To perform a cleaner analysis, let us directly examine the following row vectors:
$\mathbf{v}_{1}=[k,0,0,...,0]$ (3)
for a leaf connected to a hub with $\text{deg}(v)=k$, and
$\mathbf{v}_{2}=[1,1,1,...,1]$ (4)
for the hub, assumed to be connecting to $k$ leaves for simplicity. Both of these row vectors have a length of $k$, since extra 0’s hanging at the end does not affect cost of alignment. Moreover, suppose there is another high-degree node row vector with the same length which takes the form
$\mathbf{v}_{3}=[1+\epsilon,1+\epsilon,...,1+\epsilon].$ (5)
One might wonder whether the accumulation of small differences between each entry in the row vectors lead to a large enough difference that confuse the alignment of a high-degree node with another high-degree node versus with a low-degree node.
We assume $k\rightarrow\infty$, and $\epsilon\rightarrow 0$. Then the cost of aligning $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ is $(k-1)^{2}+(k-1)\cdot(0-1)^{2}=k(k-1)\approx k^{2}$. This is the cost of aligning a low-degree row vector with a high-degree row vector. We also compute the cost of aligning $\mathbf{v}_{2}$ with $\mathbf{v}_{3}$, which is $k\cdot(1+\epsilon-1)^{2}=k\epsilon^{2}\approx 0$. Even if we relax our assumption so that $\epsilon\rightarrow 0<C<\sqrt{k}$ rather than having $\epsilon\rightarrow 0$, the squared effect will make the alignment of high to high more reasonable.
4.3 More General Theoretical Motivations
A degree matrix ($N\times m$) encodes information more compactly than an adjacency matrix ($N\times N$), since we contain both degree distribution and each node’s local geometric details in a smaller matrix. Additionally, combining local (degrees) and global (assignment) perspectives is important because the of the Friendship Paradox. In general terms, it is the observation that friends of a person typically have more friends than that person. The Friendship Paradox shows how local observations can be completely shifted when we examine the same neighborhood globally [1, 15]. This property of graphs has made it possible for multiple definitions of “average degree” to exist on different scales, and hence led to many ways of formulating the problem to be addressed in the paradox. Intuitively, our method works well for heterogeneous graphs for the following reason. Suppose there is a graph that has nodes with degrees that are almost all unique. Note here that the phrase “almost all unique” is used to carefully address the possible scenario as follows: if we allow at most one edge between every pair of vertices on a graph $G_{example}$ with $n_{example}$ vertices and require at least degree one for all nodes, then the maximum degree is at most $n_{example}-1$ for every node, and in this case it is impossible to create a bijection between possible degrees and the nodes. With a heterogeneous graph with most degrees different, most nodes have row vectors that are relatively unique and distinguishable, compared to, for example, a complete graph.
5 Experimental Results
We first list here the sources of downloaded datasets. From COSNET [32], we downloaded the Flickr, Last.fm, and Myspace networks; from SNAP [17], we downloaded the Facebook and YouTube networks, along with the CA and PA roadmaps, and a multi-layer tissue PPI network (unweighted); from STRING [26], we downloaded five weighted PPI networks for different species that are typical subjects in biological studies; all other networks are taken from Network Repository [22]. Before we dive into results for specific graphs, we note here that all isomorphic graph alignment yielded 100% correct results.
5.1 Unweighted Graphs: DMC and Greedy DMC
5.1.1 Real World Networks